Could someone please assist me with this question:
Find the solution to this equation:
2b^2+9=0
Thank you.

Answer by stanbon(48540)

Find the solution to this equation:
2b^2+9=0
b^2 = -9/2
b = sqrt(-9/2) or b = -sqrt(-9/2)
b = 3sqrt(-1/2) or b = -3sqrt(-1/2)
b = (3/2)isqrt(2) or b = -(3/2)isqrt(2)
================
Cheers,
Stan H.